# Holomorphic Yang-Mills fields on $B$-branes

### v1

@@ 3.2.1. Generators of a derived category
Let ${\mathbf{C}}$ be an abelian category and let ${\sf E}$ be a finite set of generators of the derived category $D^{b}({\mathbf{C}}).$ Given $A$ a complex in ${\mathbf{C}}$, as usual, we denote with ${A}[l]$, with $l\in{\mathbb{Z}}$, the complex ${A}$ shifted $l$ to the left.
Given $E,\tilde{E}$ elements of the generating set ${\sf E},$ let us consider morphisms $h$ between ${E}^{\prime}:={E}[l]$ and ${\tilde{E}}^{\prime}:={\tilde{E}}[l^{\prime}]$. We denote by ${\rm Cone}(h)={E}^{\prime}[1]\oplus{\tilde{E}}^{\prime}$ the mapping cone of $h$ [11, page 154]. We define ${\sf E}^{(1)}$ the set obtained adding to ${\sf E}$ the elements of the form ${\rm Cone}(h)$. Hence, an element of ${\sf E}^{(1)}$ is a complex whose term at a position $p$ is either 00, or $E$, or a direct sum of ${E}\oplus\tilde{E},$ with $E,\tilde{E}\in{\sf E}.$
@@
Given two complexes $(A,\,\partial_{A})$ and $(B,\,\partial_{B})$ in an additive category ${\mathbf{C}}$, the complex ${\rm Hom}^{\bullet}(A,\,B)$ is defined by (see [21, page 17])
@@
Denoting by ${\rm Com}({\mathbf{C}})$ the abelian category of complexes on ${\mathbf{C}}$ and fixed the object $A,$ one has the functor $F_{A}:{\rm Com}(\mathbf{C})\to{\rm Com}(\mathbf{Ab})$ defined by
@@
Since the 00th derived functor of a left exact functor is isomorphic to the functor itself, $F_{A}=R^{0}F_{A}={\rm Ext}^{0}(A,\,.\,).$ On the other hand, ${\rm Hom}_{D({\mathbf{C}})}\big{(}{A},\,B\big{)}={\rm Ext}^{0}(A,\,{B})$ [31, Sect. 10.7], so
@@
Proof. From (3.7) and (3.9) together with the additivity of the functor ${\rm Hom}_{\mathbf{C}}(\,.\,,\,.\,),$ it follows
@@
Thus, ${\rm Hom}_{D^{b}({\mathbf{C}})}({G},\,\hat{G})=0.$
We shall deal with the particular case where the category ${\mathbf{C}}$ is a category $\mathbf{Coh}(X)$ of coherent sheaves. More precisely, let $X$ be a smooth complex $n$-dimensional variety, let $\mathbf{Qcoh}(X)$ denote the category of quasi-coherent sheaves on $X.$ One has the derived functor
$RF_{A}:D^{+}(\mathbf{Qcoh}(X))\to D^{+}(\mathbf{Ab}).$
Since ${\rm Ext}^{i}(A,\,B)=0$ for $A,B$ complexes of coherent sheaves and $i\notin[0,\,n],$ we have the functor
@@
Let ${\mathscr{G}}^{p}=\bigoplus_{pi}{\mathscr{E}}_{pi}$ be an object of $D^{b}(X)$ and $\hat{\mathscr{G}}$ is an object of this category isomorphic the complex ${\mathscr{G}}^{p}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X},$ then
${\rm Hom}_{D^{b}(X)}\big{(}{{\mathscr{G}}},\,\hat{\mathscr{G}}\big{)}=0,$
if ${\rm Hom}({\mathscr{E}}_{ip},\,\Omega^{1}_{X}({\mathscr{E}}_{jq}))=0$ for all $p,q,i,j.$
3.2.2. Generators of the derived category $D^{b}(X)$
@@
$j^{!}:\mathbf{Mod}({\mathscr{O}}_{X})\leftrightarrows\mathbf{Mod}({\mathscr{O}}_{U}):j_{!}\;\;\;\text{and}\;\;\;i^{*}:\mathbf{Mod}({\mathscr{O}}_{X})\leftrightarrows\mathbf{Mod}({\mathscr{O}}_{Z}):i_{*}.$
For every sheaf ${\mathscr{F}}$ on $X,$ the corresponding adjunction morphisms give rise to the sequence of morphisms of ${\mathscr{O}}_{X}$-modules
(3.10) $j_{!}j^{!}{\mathscr{F}}\to{\mathscr{F}}\to i_{*}i^{*}{\mathscr{F}}.$
If ${\mathscr{F}}$ is a coherent ${\mathscr{O}}_{X}$-module, then sheaf $i^{*}{\mathscr{F}}$ is coherent. Since $i$ is a proper map, $i_{*}i^{*}{\mathscr{F}}$ is a coherent ${\mathscr{O}}_{X}$-module. As $j_{!}j^{!}{\mathscr{F}}$ is a subsheaf of ${\mathscr{F}}$, it is also coherent.
In particular, when $U=X\setminus Z$, for any ${\mathscr{F}}\in\mathbf{Coh}(X)$, one has the exact sequence of ${\mathscr{O}}_{X}$-modules
(3.11) $0\to j_{!}j^{!}{\mathscr{F}}\to{\mathscr{F}}\to i_{*}i^{*}{\mathscr{F}}\to 0$
If ${\mathscr{F}}^{\bullet}$ is an object of the derived category $D^{b}(X)=D^{b}(\mathbf{Coh}(X))$, then we have the distinguished triangle
(3.12) $Rj_{!}j^{!}{\mathscr{F}}^{\bullet}\to{\mathscr{F}}^{\bullet}\to Ri_{*}i^{*}{\mathscr{F}}^{\bullet}\overset{+1}{\longrightarrow},$
where $j^{!}{\mathscr{F}}^{\bullet}\in D^{b}(\mathbf{Coh}(U))$ and $i^{*}{\mathscr{F}}^{\bullet}\in D^{b}(\mathbf{Coh}(Z)).$ Hence, as $j_{!}$ and $i_{*}$ are exact functors, one has the following proposition.
@@
If $U={\rm Spec}\,R$, where $R$ is is a Noetherian ring, the category $\mathbf{Coh}(U)$ of coherent sheaves on $U$ is equivalent to the category of finitely generated $R$-modules, with $R=\Gamma(U,\,{\mathscr{O}}_{U}).$ The functor $\Gamma(U,\,\cdot\,)$ defines the equivalence.
@@
with $S_{a}$ a finite direct sum $\oplus_{I_{a}}R.$
Denoting by $\mathbf{T}$ the triangulated subcategory of $D^{b}(\mathbf{Mod}_{f}(R))$ generated by $R,$ then the $S_{a}$ are objects of $\mathbf{T}.$ On the other hand, if we have the exact sequence $0\to S_{1}\overset{f_{1}}{\longrightarrow}S_{0}\to N\to 0$ in the category $\mathbf{Mod}_{f}(R),$ then one has the distinguished triangle
@@
Thus, $N$ belongs to the category $\mathbf{T}.$ In the case that $N$ has a resolution as (3.13), then one can consider the exact sequences
@@
An Example. Let $X$ be ${\mathbb{P}}^{1},$ $U=\{[1:x]\in{\mathbb{P}}^{1}\,|\,x\in{\mathbb{C}}\}$ and $Z=\{[0:1]\}.$ In this case, ${\mathscr{O}}_{Z}=i_{*}i^{*}{\mathscr{O}}_{X}$ is the skyscraper sheaf at the point $P=[0:1],$ and $j_{!}j^{!}{\mathscr{O}}_{X}={\mathscr{O}}_{X}(-1).$ That is, $j_{!}{\mathscr{O}}_{U}\simeq{\mathscr{O}}_{X}(-1).$
As $U={\rm Spec}\,R$ with $R$ the polynomial ring ${\mathbb{C}}[x]$, then ${\mathscr{O}}_{U}$ is the generator of the corresponding category $D^{b}(U).$ The category $D^{b}(Z)$ is generated by $i^{*}{\mathscr{O}}_{X}$, obviously. By Corollary 2), $\{{\mathscr{O}}_{X}(-1),\,{\mathscr{O}}_{Z}\}$ is a set of generators of $D^{b}(X).$
In this case, the exact sequence (3.11) is
$0\to{\mathscr{O}}_{X}(-1)\to{\mathscr{O}}_{X}\to{\mathscr{O}}_{Z}\to 0$
Therefore, the set $\{{\mathscr{O}}_{X}(-1),\,{\mathscr{O}}_{X}\}$ is also a set of generators of $D^{b}(X)$ (which is in agreement with Beilinson’s theorem). Hence, any object of $D^{b}(X)$ is isomorphic to one complex ${\mathscr{G}}^{\bullet}$ of the form (3.6), where ${\mathscr{E}}_{pi}\in\{{\mathscr{O}_{X}},\,{\mathscr{O}}_{X}(-1)\}.$
On the other hand, for ,${\mathscr{E}},\,{\mathscr{E}}^{\prime}\in\{{\mathscr{O}_{X}},\,{\mathscr{O}}_{X}(-1)\}$, one has
${\rm Hom}({\mathscr{E}},\,{\mathscr{E}}^{\prime}\otimes_{\mathscr{O}}\Omega^{1}_{X})={\rm Hom}({\mathscr{O}}_{X},\,({{\mathscr{E}}}^{\vee}\otimes{\mathscr{E}}^{\prime})\otimes_{\mathscr{O}}\Omega^{1}_{X})=H^{0}(X,\,\Omega^{1}(\epsilon)),$
where $\epsilon\in\{0,+1,-1\}$ and $\Omega^{1}(\epsilon)$ is the twisted sheaf $\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{O}}(\epsilon).$ But $H^{0}(X,\,\Omega^{1}(\epsilon))=0$ [28, page 4]. Thus, by Corollary 1 together with Proposition 5, the set of of holomorphic gauge fields over any brane on ${\mathbb{P}}^{1}$ has cardinal $\leq 1.$
Remark 5.
The family ${\sf E}=\{{\mathscr{O}}_{{\mathbb{P}}^{n}}(-n),\,{\mathscr{O}}_{{\mathbb{P}}^{n}}(-n+1),\dots,{\mathscr{O}}_{{\mathbb{P}}^{n}}\}$ generates $D^{b}({\mathbb{P}}^{n}).$ If ${\mathscr{E}},\,{\mathscr{E}}^{\prime}\in{\sf E},$ then
${\rm Hom}\big{(}{\mathscr{E}},\,\Omega^{1}_{{\mathbb{P}}^{n}}({\mathscr{E}}^{\prime})\big{)}=H^{0}\big{(}X,\,\Omega^{1}_{{\mathbb{P}}^{n}}(\epsilon)\big{)},$
with $\epsilon\in\{-n,\,-n+1,\dots,n\}.$
If $n>1,$ the cohomology group $H^{0}(X,\,\Omega^{1}_{{\mathbb{P}}^{n}}(\epsilon))$ is not necessarily zero (see [28, page 4]). Thus, Corollary 1 is not applicable in order to extend to ${\mathbb{P}}^{n}$ the above result obtained for the $B$-branes over ${\mathbb{P}}^{1}.$ To achieve such an extension it will be necessary to construct another set of generators.
3.2.3. Generators of $D^{b}({\mathbb{P}}^{n})$
We will generalize the result of the above Example. Let $X$ be a smooth a complex $n$-dimensional variety. Let $Z_{1}$ be an irreducible subvariety of $X$ of codimension $1,$ and $i_{1}:Z_{1}\hookrightarrow X$ the corresponding inclusion. We set $U_{1}:=X\setminus Z_{1}$ and $j_{1}:U_{1}\hookrightarrow X.$ Let us assume that $U_{1}$ is isomorphic to an affine variety. The exact sequence
@@ 3.2.3. Generators of $D^{b}({\mathbb{P}}^{n})$
Let us assume that there exists a tower $X=Z_{0}\supset Z_{1}\supset Z_{2}\supset\dots\supset Z_{n},$ where $Z_{k}\overset{i_{k}}{\hookrightarrow}Z_{k-1}$ is an irreducible hypersurface of $Z_{k-1}$ with $i=1,\dots,n.$ According to the above argument, ${\mathscr{M}}_{k}={\mathscr{O}}_{Z_{k-1}}(-Z_{k})$ is a sheaf supported on $U_{k}:=Z_{k-1}\setminus Z_{k}$. We set $\iota_{k}=i_{1}\circ\dots\circ i_{k}:Z_{k}\hookrightarrow X$ and $\iota_{0}:={\rm id}:Z_{0}\to X.$ For $k=1,\dots,n$ we denote ${\mathscr{L}}_{k}:=\iota_{k-1*}({\mathscr{M}}_{k}).$
We set ${\mathscr{L}}_{n+1}$ for denoting the sheaf on $X$ direct sum of the skyscraper sheaves on the points of the finite set $Z_{n}.$ Then
${\mathscr{L}}_{n+1}=\iota_{n*}\big{(}\{{\rm generators\,of}\,D^{b}(Z_{n})\}).$
We have proved the following proposition.
@@
With the above notations, the derived category $D^{b}(X)$ is generated by the set of ${\mathscr{O}}_{X}$-modules
${\sf E}=\{{\mathscr{L}}_{1},\dots,{\mathscr{L}}_{n},\,{\mathscr{L}}_{n+1}\}.$
  • When $j\neq k$ the $U_{j}\cap U_{k}=\emptyset.$ Hence, ${\mathscr{L}}_{j}$ and $\Omega^{1}({\mathscr{L}}_{k})$ have disjoint supports.; thus,
    (3.14) ${\rm Hom}_{O_{X}}\big{(}{\mathscr{L}}_{j},\,\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{L}}_{k}\big{)}=0.$
  • Since ${\mathscr{L}}_{1}$ is the invertible ${\mathscr{O}}_{X}(-Z_{1}),$ tensoring by the dual ${\mathscr{L}}_{1}^{\vee}$
    ${\rm Hom}({\mathscr{L}}_{1},\,{\mathscr{L}}_{1}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})={\rm Hom}({\mathscr{O}}_{X},\,\Omega^{1}_{X})=H^{0}(X,\,\Omega^{1}_{X})=H^{1,0}(X).$
  • In this item, when $k>1,$ we set ${\mathscr{L}}:={\mathscr{L}}_{k},$ ${\mathscr{M}}:={\mathscr{M}}_{k},$ and $\iota:=\iota_{k}$ for the inclusion $Z:=Z_{k}\hookrightarrow X;$ thus ${\mathscr{L}}=\iota_{*}{\mathscr{M}}.$ We will also delete the subscript $X$ in the sheaves ${\mathscr{O}}_{X}=:{\mathscr{O}}$ and $\Omega^{1}_{X}=:\Omega^{1}.$ By ${\mathscr{H}om}\,_{\mathscr{R}}(\,.\,,\,.\,)$ we denote the sheaf of linear homomorphisms between the corresponding ${\mathscr{R}}$-modules.
    By [22, Prop. 2.5.13] and [22, Cor. 2.3.4]
@@
    Hence,
    (3.15) ${\mathscr{H}om}\,_{\mathscr{O}}({\mathscr{L}},\,{\mathscr{L}}\otimes_{\mathscr{O}}\Omega^{1})\simeq\Gamma\big{(}X,\,\iota_{*}{\mathscr{H}om}\,_{{\mathscr{O}}_{Z}}({\mathscr{M}},\,{\mathscr{M}}\otimes_{{\mathscr{O}}_{Z}}\Omega^{1}_{Z})\big{)}=$
@@
From the above items, together with Corollary 1, Proposition 8 and Proposition 5, it follows the theorem.
@@
Let $X$ be an $n$-dimensional smooth variety and suppose that
@@
3.2.4. Holomorphic gauge fields on $B$-branes over ${\mathbb{P}}^{n}$
Let $X={\mathbb{P}}^{n}$, we define
$Z_{1}:=\{[0:x_{1}:\dots:x_{n}]\simeq{\mathbb{P}}^{n-1},\;U_{1}={\mathbb{P}}^{n}\setminus Z_{1}\simeq{\mathbb{C}}^{n}.$
Analogously, one defines $Z_{2}=\{[0:0:x_{2}:\dots:x_{n}]\}$ and $U_{2}=Z_{1}\setminus Z_{2},$ etc. Moreover, $H^{1,0}(Z_{j})=0.$ By Theorem 2, one has the following theorem.
Theorem 3.
The set of holomorphic gauge fields on any $B$-brane on ${\mathbb{P}}^{n}$ has cardinal $\leq 1.$
Remark 6.
For the $B$-brane on $X={\mathbb{P}}^{n}$ consisting only of the line bundle ${\mathscr{O}}(k),$
${\rm Hom}({\mathscr{O}}(k),\,\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{O}}(k))={\rm Hom}({\mathscr{O}},\,\Omega^{1}_{X})=H^{0}(X,\,\Omega^{1}_{X})=0.$
This fact is in agreement with the previous general result.
On the other hand, if a line bundle admits a holomorphic gauge field then its Chern class vanishes [4]. Thus, among the elements of the set $\{{\mathscr{O}}_{X}(k)\}_{k\in{\mathbb{Z}}}$, only ${\mathscr{O}}_{X}$ can support a holomorphic gauge field, and this field is precisely $\partial:{\mathscr{O}}_{X}\to\Omega^{1}_{X}.$
3.3. Flag varieties

### v2

@@ 2.3.1. Generators of a derived category
Let ${\mathbf{C}}$ be an abelian category. If $A$ a complex in ${\mathbf{C}}$, as usual, we denote by ${A}[l]$, with $l\in{\mathbb{Z}}$, the complex ${A}$ shifted $l$ to the left. Let ${\sf E}$ be a finite set of objects of ${\mathbf{C}}$ which generates the derived category $D^{b}({\mathbf{C}}).$ Given $E,\tilde{E}$ elements of the generating set ${\sf E},$ let us consider morphisms $h$ between ${E}^{\prime}:={E}[l]$ and ${\tilde{E}}^{\prime}:={\tilde{E}}[l^{\prime}]$. We denote by ${\rm Cone}(h)={E}^{\prime}[1]\oplus{\tilde{E}}^{\prime}$ the mapping cone of $h$ [9, p. 154]. We define ${\sf E}^{(1)}$ the set obtained adding to ${\sf E}$ the elements of the form ${\rm Cone}(h)$. Hence, an element of ${\sf E}^{(1)}$ is a complex whose term at a position $p$ is either 00, or $E$, or a direct sum of ${E}\oplus\tilde{E},$ with $E,\tilde{E}\in{\sf E}.$
@@
Given two complexes $(A,\,\partial_{A})$ and $(B,\,\partial_{B})$ in an additive category ${\mathbf{C}}$, the complex ${\rm Hom}^{\bullet}(A,\,B)$ is defined by (see [15, p. 17])
@@
Denoting by ${\rm Com}({\mathbf{C}})$ the abelian category of complexes on ${\mathbf{C}}$ and fixed the object $A,$ one has the functor $F_{A}:{\rm Com}(\mathbf{C})\to{\rm Com}(\mathbf{Ab}),$ where $\mathbf{Ab}$ is the category of abelian groups, defined by
@@
Since the 00th derived functor of a left exact functor is isomorphic to the functor itself, $F_{A}=R^{0}F_{A}={\rm Ext}^{0}(A,\,.\,).$ On the other hand, ${\rm Hom}_{D({\mathbf{C}})}\big{(}{A},\,B\big{)}={\rm Ext}^{0}(A,\,{B})$ [28, Sect. 10.7], so
@@
Proof. From (2.12) together with the additivity of the functor ${\rm Hom}_{\mathbf{C}}(\,.\,,\,.\,),$ it follows
@@
By (2.14), ${\rm Hom}_{D^{b}({\mathbf{C}})}({G},\,\hat{G})=0.$
We shall deal with the particular case where the category ${\mathbf{C}}$ is a category $\mathbf{Coh}(X)$ of coherent sheaves. More precisely, let $X$ be a smooth complex $n$-dimensional variety, let $\mathbf{Qcoh}(X)$ denote the category of quasi-coherent sheaves on $X.$ One has the derived functor $RF_{A}:D^{+}(\mathbf{Qcoh}(X))\to D^{+}(\mathbf{Ab}).$ Since ${\rm Ext}^{i}(A,\,B)=0$ for $A,B$ complexes of coherent sheaves and $i\notin[0,\,n],$ we have the functor
@@
Let ${\mathscr{G}}^{p}=\bigoplus_{pi}{\mathscr{E}}_{pi}$ be an object of $D^{b}(X),$ where ${\mathscr{E}}_{pi}\in\mathbf{Coh}(X),$ and $\hat{\mathscr{G}}$ is an object of $D^{b}(X)$ category isomorphic the complex $\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{G}}^{p},$ then ${\rm Hom}_{D^{b}(X)}\big{(}{{\mathscr{G}}},\,\hat{\mathscr{G}}\big{)}=0,$ if ${\rm Hom}({\mathscr{E}}_{ip},\,\Omega^{1}({\mathscr{E}}_{jq}))=0$ for all $p,q,i,j.$
Generators of the derived category $D^{b}(X)$.
@@
$\mathbf{Mod}({\mathscr{O}}_{X})\overset{j^{!}}{\to}\mathbf{Mod}({\mathscr{O}}_{U})\overset{j_{!}}{\longrightarrow}\mathbf{Mod}({\mathscr{O}}_{X})\overset{i^{*}}{\to}\mathbf{Mod}({\mathscr{O}}_{Z})\overset{i_{*}}{\longrightarrow}\mathbf{Mod}({\mathscr{O}}_{X}).$
For every sheaf ${\mathscr{F}}$ on $X,$ the corresponding adjunction morphisms give rise to the sequence of morphisms of ${\mathscr{O}}_{X}$-modules $\,j_{!}j^{!}{\mathscr{F}}\to{\mathscr{F}}\to i_{*}i^{*}{\mathscr{F}}.$ If ${\mathscr{F}}$ is a coherent ${\mathscr{O}}_{X}$-module, then sheaf $i^{*}{\mathscr{F}}$ is coherent. Since $i$ is a proper map, $i_{*}i^{*}{\mathscr{F}}$ is a coherent ${\mathscr{O}}_{X}$-module. As $j_{!}j^{!}{\mathscr{F}}$ is a subsheaf of ${\mathscr{F}}$, it is also coherent.
In particular, when $U=X\setminus Z$, for any ${\mathscr{F}}\in\mathbf{Coh}(X)$, one has the exact sequence of ${\mathscr{O}}_{X}$-modules $0\to j_{!}j^{!}{\mathscr{F}}\to{\mathscr{F}}\to i_{*}i^{*}{\mathscr{F}}\to 0$ If ${\mathscr{F}}^{\bullet}$ is an object of the derived category $D^{b}(X),$ then we have the distinguished triangle $Rj_{!}j^{!}{\mathscr{F}}^{\bullet}\to{\mathscr{F}}^{\bullet}\to Ri_{*}i^{*}{\mathscr{F}}^{\bullet}\overset{+1}{\longrightarrow},$ where $j^{!}{\mathscr{F}}^{\bullet}\in D^{b}(\mathbf{Coh}(U))$ and $i^{*}{\mathscr{F}}^{\bullet}\in D^{b}(\mathbf{Coh}(Z)).$ Hence, as $j_{!}$ and $i_{*}$ are exact functors, one has the following proposition.
@@
If $U={\rm Spec}\,R$, where $R$ is a Noetherian ring, the category $\mathbf{Coh}(U)$ of coherent sheaves on $U$ is equivalent to the category of finitely generated $R$-modules, with $R=\Gamma(U,\,{\mathscr{O}}_{U}).$ The functor $\Gamma(U,\,\cdot\,)$ defines the equivalence.
@@
with $S_{a}$ a finite direct sum $\oplus R.$
Denoting by $\mathbf{T}$ the triangulated subcategory of $D^{b}(\mathbf{Mod}_{f}(R))$ generated by $R,$ then the $S_{a}$ are objects of $\mathbf{T}.$ On the other hand, if (2.15) reduces to $0\to S_{1}\overset{f_{1}}{\longrightarrow}S_{0}\to N\to 0$ in the category $\mathbf{Mod}_{f}(R),$ then one has the distinguished triangle
@@
Thus, $N$ belongs to the category $\mathbf{T}.$ In the general case where case $N$ has a resolution as (2.15), then one can consider the exact sequences
@@
Let $X$ be a smooth a complex $n$-dimensional variety. Let $Z_{1}$ be an irreducible subvariety of $X$ of codimension $1,$ and $i_{1}:Z_{1}\hookrightarrow X$ the corresponding inclusion. We set $U_{1}:=X\setminus Z_{1}$ and $j_{1}:U_{1}\hookrightarrow X.$ Let us assume that $U_{1}$ is isomorphic to an affine variety. The exact sequence
@@
Let us assume that there exists a tower $X=Z_{0}\supset Z_{1}\supset Z_{2}\supset\dots\supset Z_{n},$ where $Z_{k}\overset{i_{k}}{\hookrightarrow}Z_{k-1}$ is an irreducible hypersurface of $Z_{k-1}$ with $i=1,\dots,n.$ According to the above argument, ${\mathscr{M}}_{k}={\mathscr{O}}_{Z_{k-1}}(-Z_{k})$ is a sheaf supported on $U_{k}:=Z_{k-1}\setminus Z_{k}$. We set $\iota_{k}=i_{1}\circ\dots\circ i_{k}:Z_{k}\hookrightarrow X$ and $\iota_{0}:={\rm id}:Z_{0}\to X.$ For $k=1,\dots,n$ we denote ${\mathscr{L}}_{k}:=\iota_{k-1*}({\mathscr{M}}_{k}).$ We set ${\mathscr{L}}_{n+1}$ for denoting the sheaf on $X$ direct sum of the skyscraper sheaves on the points of the finite set $Z_{n}.$ Then ${\mathscr{L}}_{n+1}=\iota_{n*}\big{(}\{{\rm generators\,of}\,D^{b}(Z_{n})\}).$ We have proved the following proposition.
@@
With the above notations, the derived category $D^{b}(X)$ is generated by the set of ${\mathscr{O}}_{X}$-modules $\,{\sf E}=\{{\mathscr{L}}_{1},\dots,{\mathscr{L}}_{n},\,{\mathscr{L}}_{n+1}\}.$
Remark 4.
When $j\neq k$ the intersection $U_{j}\cap U_{k}=\emptyset.$ Hence, ${\mathscr{L}}_{j}$ and $\Omega^{1}({\mathscr{L}}_{k})$ have disjoint supports; thus,
(2.16) ${\rm Hom}\big{(}{\mathscr{L}}_{j},\,\Omega^{1}_{X}\otimes_{{\mathscr{O}}_{X}}{\mathscr{L}}_{k}\big{)}=0.$
Lemma 2.
With the notation introduced above, ${\rm Hom}({\mathscr{L}}_{k},\,{\mathscr{L}}_{k}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})=H^{1,0}(Z_{k}),$ for $k=1,\dots,n.$
Proof. We set in this proof ${\mathscr{L}}:={\mathscr{L}}_{k},$ ${\mathscr{M}}:={\mathscr{M}}_{k},$ and $\iota:=\iota_{k-1}$ for the inclusion $Z:=Z_{k-1}\hookrightarrow X;$ thus ${\mathscr{L}}=\iota_{*}{\mathscr{M}}.$ We will also delete the subscript $X$ in the sheaves ${\mathscr{O}}_{X}=:{\mathscr{O}}$ and $\Omega^{1}_{X}=:\Omega^{1}.$ By ${\mathscr{H}om}\,_{\mathscr{R}}(\,.\,,\,.\,)$ we denote the sheaf of linear homomorphisms between the corresponding ${\mathscr{R}}$-modules.
By [16, Prop. 2.5.13] and [16, Cor. 2.3.4]
@@
By [16, Prop. 2.5.13] and [16, Cor. 2.3.4]
Hence, as ${\mathscr{M}}$ is an invertible ${\mathscr{O}}_{Z}$-module,
${\rm Hom}({\mathscr{L}},\,{\mathscr{L}}\otimes_{\mathscr{O}}\Omega^{1})\simeq\Gamma\big{(}X,\,\iota_{*}{\mathscr{H}om}\,_{{\mathscr{O}}_{Z}}({\mathscr{M}},\,{\mathscr{M}}\otimes_{{\mathscr{O}}_{Z}}\Omega^{1}_{Z})\big{)}=$
@@
Let $X$ be an $n$-dimensional smooth variety and suppose that:
@@
Proof. From Lemma 2 together with Remark 4, it follows that ${\rm Hom}({\mathscr{L}}_{j},\,\Omega^{1}({\mathscr{L}}_{k}))=0,$ for all $j,k.$ Given a brane ${\mathscr{F}}^{\bullet}$ over $X,$ from Corollary 1 and Proposition 4, we deduce that ${\rm Hom}_{D^{b}(X)}({\mathscr{F}}^{\bullet},\,\Omega^{1}({\mathscr{F}}^{\bullet}))=0.$ Now the theorem is consequence of Proposition 1. ∎
Using this theorem when $X$ is the projective space, it has been proven in [27] that the set of holomorphic gauge fields on any $B$-brane on ${\mathbb{P}}^{n}$ has cardinal $\leq 1.$
2.3.2. Flag varieties